My son need help with his math homework,we have tried to work it since he been home from school the problem is;
Dennis mowed his next door neighbor's lawn for a handful of dimes and nickels, 80 coins in all. Upon completing the job he countesd out the coins and it came to to $6.60. how many of each coin did he earn?
Have to use a linear system to solve

Answer by ankor@dixie-net.com(22740)

Dennis mowed his next door neighbor's lawn for a handful of dimes and nickels, 80 coins in all.
Upon completing the job he counted out the coins and it came to to $6.60.
how many of each coin did he earn?
:
Let n = no. of nickels; let d = no. of dimes
:
The no. of coins equation:
n + d = 80
or
n = (80-d), use this for substitution in the value equation
:
The total value equation
.05n + .10d = 6.60
Replace n with (80-d)
.05(80-d) + .10d = 6.60
4 - .05d + .10d = 6.60
-.05d + .10d = 6.60 - 4
.05d = 2.60
d = 2.6/.05
d = 52 dimes
then
80 - 52 = 28 nickels
:
:
See if that come out right
.05(28) + .10(52) =
1.40 + 5.20 = 6.60 confirms our solution of 28 nickels and 52 dimes